# assumption[theorem]Assumption

### v1

@@ 3.4 Boundary case: two-term bulk expansion
Proof 3.17.
Insert $\Lambda=(\pi-C)/\epsilon$ into the two-term Weyl expansion and use Lemma 2.7.
@@
as claimed.
4 Heat-trace and zeta correspondences in the edge variable
4.1 Heat trace transfer
Lemma 4.1 (Heat trace convergence and small-$t$ asymptotics).
On a smooth compact closed $(M^{d},g)$ the heat trace
$\Theta_{\Delta}(t):=\sum_{n}m_{n}e^{-t\lambda_{n}}$ (4.1)
converges for every $t>0$ and admits, as $t\downarrow 0$, the Seeley–DeWitt expansion
$\Theta_{\Delta}(t)\sim(4\pi t)^{-d/2}\sum_{j=0}^{\infty}a_{j}\,t^{j/2},$ (4.2)
where $a_{0}=\mathrm{Vol}(M)$ and $a_{j}$ are integrals of local curvature invariants; [3] Define the spectral zeta function by
$\zeta_{\Delta}(u):=\sum_{n}m_{n}\,\lambda_{n}^{-u}\qquad(\Re u>d/2).$ (4.3)
Consequently, $\zeta_{\Delta}$ admits a meromorphic continuation to $\mathbb{C}$ with at most simple poles located at $u=\frac{d-j}{2}$ for $j=0,1,2,\dots$.
Lemma 4.2 (Bounds for heat coefficients).
For each $j\geq 0$ there exists a universal polynomial $P_{j}$ in the curvature tensor and its covariant derivatives up to order $j$ such that
$a_{j}=\int_{M}P_{j}(\mathrm{Rm},\nabla\mathrm{Rm},\dots,\nabla^{j}\mathrm{Rm})\,d\mathrm{vol}_{g}.$ (4.4)
In particular, on a fixed compact $(M,g)$ one has $|a_{j}|\leq C_{j}(M,g)$.
Proof 4.3.
See, e.g., [10, Ch. 4] and [11].
@@ 4.1 Heat trace transfer
Assume the purely affine encoding $C_{n}=\pi-\epsilon\lambda_{n}$ (i.e. $\delta\equiv 0$). With $\Theta_{\Delta}$ and $\zeta_{\Delta}$ as above, define
@@
Hence the pole structure of $\zeta_{\mathrm{edge}}$ is identical to $\zeta_{\Delta}$.
Proof 4.5.
Since $\pi-C_{n}=\epsilon\lambda_{n}$ for all $n$, we have $H_{\mathrm{edge}}(s)=\sum_{n}m_{n}e^{-s(\pi-C_{n})}=\sum_{n}m_{n}e^{-\epsilon s\,\lambda_{n}}=\Theta_{\Delta}(\epsilon s)$, and $\zeta_{\mathrm{edge}}(u)=\sum_{n}m_{n}(\pi-C_{n})^{-u}=\epsilon^{-u}\sum_{n}m_{n}\lambda_{n}^{-u}=\epsilon^{-u}\zeta_{\Delta}(u)$.
Remark 4.6 (Effect of non-affine perturbations).
If $C_{n}=\pi-\epsilon\lambda_{n}+\delta(\lambda_{n})$ with $\delta(\lambda)=o(\lambda)$, then
$H_{\mathrm{edge}}(s)=\sum_{n}m_{n}e^{-\epsilon s\,\lambda_{n}}\,e^{\,s\,\delta(\lambda_{n})},\qquad\zeta_{\mathrm{edge}}(u)=\sum_{n}m_{n}\big(\epsilon\lambda_{n}-\delta(\lambda_{n})\big)^{-u}.$ (4.7)
Thus the exact identities of Proposition 4.4 are lost, although for asymptotic questions one recovers the leading behavior under the hypotheses of Section 6.
4.2 Spectral zeta transfer
Corollary 4.7.
For $\Re u>d/2$, $\zeta_{\mathrm{edge}}(u)=\epsilon^{-u}\,\zeta_{\Delta}(u)$. Consequently, the meromorphic continuation and residues of $\zeta_{\mathrm{edge}}$ are inherited from $\zeta_{\Delta}$ under the trivial scaling by $\epsilon^{-u}$.
Proof 4.8.
Immediate from Proposition 4.4.
@@ 5.1 Polynomial-type encodings and inverse asymptotics
Theorem 5.1 (Uniqueness among polynomial-type encodings).
@@
Proof 5.2.
We have $\phi(\lambda)=a-g(\lambda)\sim b\,\lambda^{k}L(\lambda)$, so $\phi\in\mathrm{RV}_{k}$ and is ultimately increasing. By inversion theory for regularly varying functions (de Bruijn conjugacy; [9, §1.5, Thm. 1.5.12]), there exists an asymptotic inverse $\Lambda\in\mathrm{RV}_{1/k}$ with
@@
Let $F(x):=N_{\mu_{g}}(a-x)$. By the monotone density theorem ([9, §1.7, Thm. 1.7.2]), $F^{\prime}(x)\in\mathrm{RV}_{\,d/(2k)-1}$. Since $\rho_{\mu_{g}}(C)=-\frac{d}{dC}N_{\mu_{g}}(C)=\frac{d}{dx}F(x)$ with x=aCx=a-C, the edge density exponent is d/(2k)1d/(2k)-1. Matching this with (d2)/2(d-2)/2 forces d/(2k)1=(d2)/2d/(2k)-1=(d-2)/2, i.e. $k=1$. Slowly varying factors $L,L^{\ast}$ affect the prefactor but not the exponent.
@@
Proof $5.$
Immediate from Theorem 5.1 and the monotone density theorem [9, §1.7].
5.2 Edge exponents and the necessity of $k=1$
Proposition 5.6 (Edge-variable exponents under $g(\lambda)\sim a-b\lambda^{k}L(\lambda)$).
Let $g(\lambda)\sim a-b\,\lambda^{k}L(\lambda)$ as $\lambda\to\infty$ with $a\in\mathbb{R}$, $b>0$, $k>0$, and $L\in\mathrm{RV}_{0}$, and assume $g$ is ultimately strictly decreasing (by strict monotonicity of $g$; cf. Lemma 2.4 for the affine case). Put $x:=a-C\downarrow 0$ and let $\mu_{g}$ be the pushforward of $\mu_{\Delta}$ by $g$. Then
$N_{\mu_{g}}(a-x)\in\mathrm{RV}_{\,d/(2k)},\qquad\rho_{\mu_{g}}(a-x)\in\mathrm{RV}_{\,d/(2k)-1}.$ (5.6)
In particular, the edge-variable counting exponent equals $d/(2k)$ and the edge-variable density exponent equals d/(2k)1d/(2k)-1.
Proof 5.7.
This is an immediate consequence of Theorem 5.1: writing $\phi(\lambda)=a-g(\lambda)\in\mathrm{RV}_{k}$ and letting $\Lambda$ be an asymptotic inverse, one has $N_{\mu_{g}}(a-x)=N_{\Delta}(\Lambda(x))\sim\gamma_{d}\,\Lambda(x)^{d/2}\in\mathrm{RV}_{d/(2k)}$, and then the monotone density theorem yields the claim for $\rho_{\mu_{g}}$.
Corollary 5.8 (Necessity of $k=1$ to match the geometric edge-variable exponent).
The geometric edge-variable density exponent is (d2)/2(d-2)/2. Matching this with the value d/(2k)1d/(2k)-1 from Proposition 5.6 forces $k=1$. Hence any polynomial-type encoding preserving the edge-variable exponent must be asymptotically affine:
$g(\lambda)=a-b\lambda\,L(\lambda),\qquad L\in\mathrm{RV}_{0}.$ (5.7)
Slow variation $L$ cannot alter the exponent (only the slowly varying prefactor).
Remark 5.9 (What happens for $k\neq 1$).
If $k>1$ the edge-variable density exponent d/(2k)1d/(2k)-1 is strictly smaller than (d2)/2(d-2)/2, yielding an artificially thinner edge density; if $k<1$ it is strictly larger, yielding an artificially thicker edge density. In both cases the mismatch cannot be corrected by any slowly varying factor $L$.
@@ 5.3 Slow variation and prefactors
and the monotone density theorem yields the claim for $\rho_{\mu_{g}}$.
@@
When $k=1$ (the asymptotically affine case required by Corollary 5.8):
@@ 5.4 Uniqueness of the affine scaling from the bulk exponent
Theorem 5.16 (Bulk exponent singles out index $1$ in full $O$-regular variation).
@@
Proof 5.17.
(i)–(ii) follow from standard inversion properties of $O$-regularly varying functions and bounding the inverse via the upper/lower Matuszewska indices [9, §§2.1–2.4]. Composing with Weyl’s law $N_{\Delta}(\Lambda)\sim\gamma_{d}\,\Lambda^{d/2}$ transfers the index interval. (iii) A single sharp power exponent for $N_{\mu_{g}}$ (or for $\rho_{\mu_{g}}$ via the monotone density theorem) forces the upper and lower indices of $\Lambda$ to coincide, hence those of $\phi$ coincide and equal $1$, i.e. $\phi\in\mathrm{RV}_{1}$. Then $\phi(\lambda)=b\lambda L(\lambda)$ with $L\in\mathrm{RV}_{0}$ by Karamata’s representation [9, Thm. 1.3.1], which yields the stated form of $g$.
@@
If $\phi$ is faster than any power (e.g. $\phi(\lambda)=e^{\lambda}$), then $\Lambda(x)\asymp\log(1/x)$ and $N_{\mu_{g}}(a-x)\asymp(\log(1/x))^{d/2}$, $\rho_{\mu_{g}}(a-x)\asymp x^{-1}(\log(1/x))^{d/2-1}$, which is not a power law. If $\phi(\lambda)=\lambda^{k}L(\lambda)$ with $k\neq 1$, Proposition 5.6 shows the exponent mismatch $d/(2k)-1\neq(d-2)/2$.
@@ 6.1 Setup and inversion error
Combining the four steps completes the proof.
@@ 6.2 Preservation of edge-variable exponents
Since $\rho_{\mathrm{bulk}}(C)=-\frac{d}{dC}N_{\mu_{C}}(C)=\frac{d}{dy}F(y)$ with $y=\pi-C$, the stated asymptotics for $\rho_{\mathrm{bulk}}$ follow.
@@
Proof $6.$
Integrating $|\delta^{\prime}(\lambda)|\leq\ell(\lambda)/\lambda$ yields $\delta(\lambda)=o(\lambda)$ with explicit control; then apply Lemma 6.1 and the proof of Proposition 6.3, using the uniform convergence theorem for slowly varying functions [9, §1.8].
@@ 6.3 Examples of admissible perturbations
Proof 6.8.
Items (a), (b), (c), (f) follow from Proposition 6.3 with the displayed choices of $\eta\in\mathrm{RV}_{0}$ (uniform convergence for slowly varying functions supplies the replacement $\eta(\Lambda)\rightsquigarrow\eta(y/\epsilon)$). Item (e) follows from Lemma 6.5 with $\ell(\lambda)=\beta(\log(e\lambda))^{\beta-1}$. Item (d) is obtained by a one–step inversion of $y=\epsilon\Lambda-\delta(\Lambda)$ with $\delta=O(1)$, which yields $\Lambda=y/\epsilon+O(1)$, and then by applying Weyl’s law and Karamata’s monotone density theorem to pass to $N_{\mu_{C}}$ and $\rho_{\mathrm{bulk}}$.
@@
7 Model geometries and multiplicities

### v2

@@ 3.3 Boundary case: two-term bulk expansion
Although Assumption 2.1 requires compact manifolds without boundary, the following extension to the boundary case is standard [21, 7, 8] and recorded for completeness. When $M$ has smooth boundary, Weyl’s law acquires a second term proportional to the boundary area. The following proposition shows that this two-term structure transfers intact to the edge variable, with both terms rescaled by the appropriate power of $\epsilon$.
@@
Proof 3.11.
Insert $\Lambda=(\pi-C)/\epsilon$ into the two-term Weyl expansion and use Lemma 2.7. ∎
@@
as claimed. ∎
3.4 Heat-trace and zeta transfer
The affine encoding transfers the heat trace and spectral zeta to the edge variable by the same rescaling identity that underlies the Weyl transfer above.
@@ 3.4 Heat-trace and zeta transfer
Under the purely affine encoding $C_{n}=\pi-\epsilon\lambda_{n}$, define
@@
Consequently, all Seeley–DeWitt heat invariants [24, 11, 10] and the pole structure of $\zeta_{\mathrm{edge}}$ are preserved up to explicit $\epsilon$-scalings; for constant-curvature spaces this gives $a_{2m}^{\mathrm{edge}}=\epsilon^{m-d/2}a_{2m}$ (see Section 6.1). Non-affine perturbations $C_{n}=\pi-\epsilon\lambda_{n}+\delta(\lambda_{n})$ with $\delta=o(\lambda)$ break these exact identities; asymptotic recovery is treated in Section 5.
Proof 3.15.
Since $\pi-C_{n}=\epsilon\lambda_{n}$ for all $n$, both identities follow by direct substitution: $H_{\mathrm{edge}}(s)=\sum_{n}m_{n}e^{-\epsilon s\lambda_{n}}=\Theta_{\Delta}(\epsilon s)$ and $\zeta_{\mathrm{edge}}(u)=\epsilon^{-u}\sum_{n}m_{n}\lambda_{n}^{-u}=\epsilon^{-u}\zeta_{\Delta}(u)$. ∎
@@ 4 Uniqueness of the affine encoding
This section proves the two uniqueness results stated in Theorem 1.1. The main result is Theorem 4.12 in Section 4.4: the bulk power-law constraint forces $\phi\in\mathrm{RV}_{1}$ within the full O-regularly varying class. Section 4.1 recovers the polynomial case $k=1$ as a direct specialization via de Bruijn’s inversion theorem [9, 26]; the remaining subsections analyze prefactors and slowly varying corrections.
@@ 4.1 Polynomial-type encodings and inverse asymptotics
The following specializes the rigidity to the polynomial class $\mathrm{RV}_{k}$; the self-contained proof via de Bruijn inversion makes the exponent formula explicit. The general O-RV statement (Theorem 4.12) is established in Section 4.4 below.
Theorem 4.1 (Polynomial specialization).
@@
Proof 4.2.
We have $\phi(\lambda)=a-g(\lambda)\sim b\,\lambda^{k}L(\lambda)$, so $\phi\in\mathrm{RV}_{k}$ and is ultimately increasing. By inversion theory for regularly varying functions (de Bruijn conjugacy; [9, §1.5, Thm. 1.5.12]; see also [26]), there exists an asymptotic inverse $\Lambda\in\mathrm{RV}_{1/k}$ with
@@
Let $F(x):=N_{\mu_{g}}(a-x)$. By the monotone density theorem ([9, §1.7, Thm. 1.7.2]), $F^{\prime}(x)\in\mathrm{RV}_{\,d/(2k)-1}$. Since $\rho_{\mu_{g}}(C)=-\frac{d}{dC}N_{\mu_{g}}(C)=\frac{d}{dx}F(x)$ with x=aCx=a-C, the edge density exponent is d/(2k)1d/(2k)-1. The nontrivial content is the inverse transfer $\phi\in\mathrm{RV}_{k}\Rightarrow\Lambda\in\mathrm{RV}_{1/k}$ via de Bruijn conjugacy (4.2) and the monotone density theorem; the exponent matching is then an unavoidable consequence: d/(2k)1=(d2)/2d/(2k)-1=(d-2)/2 forces $k=1$. Slowly varying factors $L,L^{\ast}$ affect the prefactor but not the exponent. ∎
@@
Proof 4.5.
Immediate from Theorem 4.1 and the monotone density theorem [9, §1.7]. ∎
4.2 Recovering $k$ from edge-variable data
By Theorem 4.1, the exponent formula $N_{\mu_{g}}(a-x)\in\mathrm{RV}_{d/(2k)}$ immediately yields a practical estimator for $k$.
@@ 4.3 Slow variation and prefactors
While $k=1$ is forced by the exponent constraint, the slowly varying factor $L\in\mathrm{RV}_{0}$ in the affine encoding $g(\lambda)=a-b\lambda L(\lambda)$ can still affect the prefactor via the de Bruijn conjugate $L^{\ast}$. This subsection quantifies that effect and shows the prefactor reduces to the canonical form when $L\to 1$.
@@
and the monotone density theorem yields the claim for $\rho_{\mu_{g}}$. ∎
@@
When $k=1$ (the asymptotically affine case required by Theorem 4.1):
@@ 4.4 Uniqueness of the affine scaling from the bulk exponent
The following is the main rigidity theorem of this section; Theorem 4.1 above is its polynomial specialization.
Theorem 4.12 (O-RV rigidity: main theorem).
@@
Proof 4.13.
(i)–(ii) follow from standard inversion properties of $O$-regularly varying functions and bounding the inverse via the upper/lower Matuszewska indices [9, §§2.1–2.4]. Composing with Weyl’s law $N_{\Delta}(\Lambda)\sim\gamma_{d}\,\Lambda^{d/2}$ transfers the index interval. (iii) A single sharp power exponent for $N_{\mu_{g}}$ (or for $\rho_{\mu_{g}}$ via the monotone density theorem) forces the upper and lower indices of $\Lambda$ to coincide, hence those of $\phi$ coincide and equal $1$, i.e. $\phi\in\mathrm{RV}_{1}$. Then $\phi(\lambda)=b\lambda L(\lambda)$ with $L\in\mathrm{RV}_{0}$ by Karamata’s representation [9, Thm. 1.3.1], which yields the stated form of $g$. ∎
@@
If $\phi$ is faster than any power (e.g. $\phi(\lambda)=e^{\lambda}$), then $\Lambda(x)\asymp\log(1/x)$ and $N_{\mu_{g}}(a-x)\asymp(\log(1/x))^{d/2}$, $\rho_{\mu_{g}}(a-x)\asymp x^{-1}(\log(1/x))^{d/2-1}$, which is not a power law. If $\phi(\lambda)=\lambda^{k}L(\lambda)$ with $k\neq 1$, Theorem 4.1 shows the exponent mismatch $d/(2k)-1\neq(d-2)/2$.
4.5 Asymptotic spectral classes and encoding-induced morphisms
Definition 4.15 (Asymptotic spectral equivalence class).
Let $\mu_{1},\mu_{2}$ be positive spectral measures with ultimately monotone counting functions $N_{1},N_{2}$. For $\beta>0$ write
$\mu_{1}\sim_{\mathrm{as},\beta}\mu_{2}\quad\Longleftrightarrow\quad N_{1}\in\mathrm{RV}_{\beta}\ \text{and}\ N_{2}\in\mathrm{RV}_{\beta}.$ (4.22)
The corresponding class is denoted $\mu]_{\mathrm{as},\beta}$. Whenever $N_{\mu}\in\mathrm{RV}_{\beta}$, define the asymptotic spectral dimension
$d_{\mathrm{as}}(\mu):=2\beta.$ (4.23)
Proposition 4.16 (Encoding-induced dimension scaling).
Assume $N_{\Delta}(\Lambda)\sim\gamma_{d}\Lambda^{d/2}$ and let $g(\lambda)=a-\phi(\lambda)$ be ultimately strictly decreasing with $\phi\in\mathrm{RV}_{k}$, $k>0$. Then
$N_{\mu_{g}}(a-x)\in\mathrm{RV}_{d/(2k)},$ (4.24)
so $g$ induces a map between asymptotic classes
$T_{g}:\ [\mu_{\Delta}]_{\mathrm{as},\,d/2}\longrightarrow[\mu_{g}]_{\mathrm{as},\,d/(2k)},$ (4.25)
and therefore acts on asymptotic spectral dimension by
$d_{\mathrm{as}}\longmapsto\frac{1}{k}\,d_{\mathrm{as}}.$ (4.26)
Proof 4.17.
This is exactly Theorem 4.1: if $\phi\in\mathrm{RV}_{k}$ then $N_{\mu_{g}}(a-x)=N_{\Delta}(\Lambda(x))\in\mathrm{RV}_{d/(2k)}$, where $\Lambda$ is an asymptotic inverse of $\phi$. ∎
Corollary 4.18 (Dimension-preserving morphisms).
Under the hypotheses of Theorem 4.12:
  1. (i)
    $g(\lambda)=a-\phi(\lambda)$ preserves asymptotic spectral dimension if and only if $N_{\mu_{g}}(a-x)\in\mathrm{RV}_{d/2}$ (equivalently $\rho_{\mu_{g}}\in\mathrm{RV}_{(d-2)/2}$);
  2. (ii)
    if a genuine edge power law holds, namely
    then necessarily $\phi\in\mathrm{RV}_{1}$;
  3. (iii)
    conversely, $\phi\in\mathrm{RV}_{1}$ implies $N_{\mu_{g}}(a-x)\in\mathrm{RV}_{d/2}$ and hence dimension preservation.
If in addition $\phi(\lambda)=\epsilon\lambda L(\lambda)$ with $L(\lambda)\to 1$, then
$N_{\mu_{g}}(a-x)\sim\gamma_{d}\,\epsilon^{-d/2}x^{d/2},$ (4.28)
and the normalization is asymptotically strict affine: $g(\lambda)=a-\epsilon\lambda+o(\lambda)$.
Proof 4.19.
The first statement is the definition of $d_{\mathrm{as}}=2\beta$. The implication in (ii) is Theorem 4.12(iii). For (iii), if $\phi\in\mathrm{RV}_{1}$, then its asymptotic inverse satisfies $\Lambda\in\mathrm{RV}_{1}$, and therefore $N_{\mu_{g}}(a-x)=N_{\Delta}(\Lambda(x))\in\mathrm{RV}_{d/2}$ by Weyl’s law and composition of regularly varying functions. The normalization statement follows from Corollary 4.11. ∎
Remark 4.20 (Hierarchy of preservation).
  1. (i)
    preserving Weyl exponent (equivalently asymptotic dimension) is exactly $\phi\in\mathrm{RV}_{1}$;
  2. (ii)
    preserving the normalized Weyl constant requires the first-order normalization $\phi(\lambda)=\epsilon\lambda(1+o(1))$;
  3. (iii)
    this is the strict affine regime at leading order.
@@ 5 Stability under perturbations
The uniqueness results of Section 4 concern exact encodings. Here we show that the geometric bulk exponents are stable when the affine encoding is perturbed by $\delta(\lambda)=o(\lambda)$: the counting function and bulk density retain the same power-law exponents, with error rates controlled by the slowly varying factor $\eta=\delta/\lambda\in\mathrm{RV}_{0}$ [9]. This stability is relevant whenever the encoding is known only approximately, for instance when $\epsilon$ is estimated from data or when the encoding is subject to a systematic drift.
@@ 5.1 Setup and inversion error
The starting point is an explicit inversion estimate: given the perturbed encoding $C(\lambda)=\pi-\epsilon\lambda+\delta(\lambda)$, we bound the deviation of the inverse $\Lambda(C)$ from the unperturbed value $y/\epsilon$.
@@
Combining the four steps completes the proof. ∎
@@ 5.2 Preservation of edge-variable exponents
Using the inversion bound from Lemma 5.1, we now state the main stability result: all three quantities $\Lambda(C)$, $N_{\mu_{C}}(C)$, and $\rho_{\mathrm{bulk}}(C)$ have the same asymptotics as in the unperturbed case, up to a multiplicative factor $1+O(\eta(y/\epsilon))$.
@@
Since $\rho_{\mathrm{bulk}}(C)=-\frac{d}{dC}N_{\mu_{C}}(C)=\frac{d}{dy}F(y)$ with $y=\pi-C$, the stated asymptotics for $\rho_{\mathrm{bulk}}$ follow. ∎
@@
Proof 5.6.
Integrating $|\delta^{\prime}(\lambda)|\leq\ell(\lambda)/\lambda$ yields $\delta(\lambda)=o(\lambda)$ with explicit control; then apply Lemma 5.1 and the proof of Proposition 5.3, using the uniform convergence theorem for slowly varying functions [9, §1.8]. ∎
@@ 5.3 Examples of admissible perturbations
We illustrate the stability result with explicit families of perturbations, computing the slowly varying factor $\eta$ and the resulting bulk error profile in each case.
@@
Proof 5.8.
Items (a), (b), (c), (f) follow from Proposition 5.3 with the displayed choices of $\eta\in\mathrm{RV}_{0}$ (uniform convergence for slowly varying functions supplies the replacement $\eta(\Lambda)\rightsquigarrow\eta(y/\epsilon)$). Item (e) follows from Lemma 5.5 with $\ell(\lambda)=\beta(\log(e\lambda))^{\beta-1}$. Item (d) is obtained by a one–step inversion of $y=\epsilon\Lambda-\delta(\Lambda)$ with $\delta=O(1)$, which yields $\Lambda=y/\epsilon+O(1)$, and then by applying Weyl’s law and Karamata’s monotone density theorem to pass to $N_{\mu_{C}}$ and $\rho_{\mathrm{bulk}}$. ∎
@@
6 Geometric models and Kreĭn realization
We specialize the general theory to explicit geometric models. Section 6.1 computes the exact scaling of Seeley–DeWitt heat coefficients [24, 11, 10] for constant-curvature spaces, confirming the alignment $a_{2m}^{\mathrm{edge}}=\epsilon^{m-d/2}a_{2m}$. Section 6.2 addresses the realization of the full spectral measure with multiplicities: a Kreĭn string [12, 14] realizes any such measure uniquely, while no classical smooth Sturm–Liouville operator can do so for $d>1$ [17, 13]. Section 6.3 analyzes spherical eigenvalue clustering and shows it averages out in the bulk without affecting the Weyl exponent.